# DISTINQT: A Distributed Privacy Aware Learning Framework for QoS Prediction for Future Mobile and Wireless Networks

### v1

@@ IV-A Configuration Setup
Network Environment: The discrete event network simulator (NS3) [44] has been used for the performance evaluation of the DISTINQT framework. The simulated topologies and the evaluation methodology chosen for this study are in line with the 3GPP’s guidelines [45], [46]. The duration of the simulation is equal to 280 minutes, while the logging frequency of the data is equal to $\tau_{e}=200$ milliseconds (ms) for all NETs in our study. Two types of UEs are considered in the simulation, namely ToD-UE and NonToD-UE vehicles, one BS and one MEC server. According to [47], the periodic traffic of the ToD-UEs is set to 20 Mbps for the uplink and 500 Kbps for the downlink. On the other hand, the NonToD-UEs generate periodic background traffic load that interferes with that from the ToD-UE vehicles and will impact on the latter’s QoS. The number of NonToD-UEs varies over time and is randomly selected in the range of $[0,100]$. Each NonToD-UE generates traffic of 4 Mbps for the uplink and 8 Kbps for the downlink, considered to be constant throughout the experiments. The mobility of all UEs was generated using the Simulation of Urban Mobility tool (SUMO) [48]. The ToD-UEs are moving continuously with a variable speed in the range of $(0,14]$ m/s corresponding to $(0,50]$ km/h, while the NonToD-UEs follow randomly generated trajectories inside the simulated topology, as provided by SUMO. We consider 10 simulated experiments, differing in the total background traffic load (depending on the number of NonToD-UEs) and the mobility patterns of the UEs. Table II summarizes the NS-3 parameters used in the simulated scenario.
@@
The maximum number of epochs for the learning phase was set to 1000. Additionally, in order to avoid overfitting and potential degradation in the performance of the model, an early stopping function is applied, which monitors the Mean Squared Error (MSE) of the validation set [49]. As a result, the learning process stops, when the MSE score does not improve or starts deteriorating.
@@ V-A DISTINQT vs Centralized
We compare the performance of our DISTINQT framework with a centralized LSTM-based approach, introduced in our previous work [51]. Table V, presents the evaluation results as average values of the MAE score for both input feature configurations and for all evaluation scenarios (50 in total). Two types of evaluation metrics are considered: (a) The ”overall MAE” score, representing the average MAE score over all prediction horizon timesteps, and (b) the ”20th Second MAE” score, representing the average MAE score of the last prediction horizon timestep only. The results under both the centralized, [51], and our proposed DISTINQT approaches are statistically very similar, with the DISTINQT showcasing a marginally better performance of 7.75% on average for both configurations and types of scores considered. Also c2 improves upon c1 on average by 5.65% for both scores, highlighting the importance of the additional input features of the MEC Server NET. MAE scores capture the overall ability of the model to yield a relatively low prediction error under the DISTINQT and the centralized approaches, considering typical uplink throughput values of the ToD applications (10-50 Mbps), as indicated by the EU funded 5GCroCo project [17].
@@ VI Conclusions
This paper introduces and demonstrate the effectiveness of DISTINQT, a novel privacy-aware distributed learning framework for QoS prediction in the presence of heterogeneous data types and model architectures. DISTINQT is based on sequence-to-sequence autoencoders, distributing various NN-Encoders across the network and contributing to data privacy preservation. Our framework is evaluated in the Tele-Operated Driving use case, where evaluation results show that it can reach and even surpass the performance of its centralized counterpart. This is especially important for such safety critical use cases that require immediate and accurate decision making without compromising any sensitive information. Finally, DISTINQT achieves an average performance improvement of up to 65%, compared to six state-of-the-art centralized baseline solutions. Our future work will focus on finding a balance between the communication overhead that is introduced in the learning process and the robustness and generalization capabilities of the model.

### v2

@@ IV-A Configuration Setup
Network Environment: The discrete event network simulator (NS3) [47] has been used for the performance evaluation of the DISTINQT framework. The simulated topologies and the evaluation methodology chosen for this study are in line with the 3GPP’s guidelines [48], [49]. More specifically, the simulated topology chosen for this study is an Urban Macro deployment for connected cars, known as Manhattan Grid. The duration of the simulation is equal to 280 minutes, while the logging frequency of the data is equal to $\tau_{e}=200$ milliseconds (ms) for all NETs in our study. Two types of UEs are considered in the simulation, namely ToD-UE and NonToD-UE vehicles, one BS and one MEC server. According to [50], the periodic traffic of the ToD-UEs is set to 20 Mbps for the uplink and 500 Kbps for the downlink. On the other hand, the NonToD-UEs generate periodic background traffic load that interferes with that from the ToD-UE vehicles and will impact on the latter’s QoS. The number of NonToD-UEs varies over time and is randomly selected in the range of $[0,100]$. Each NonToD-UE generates traffic of 4 Mbps for the uplink and 8 Kbps for the downlink, considered to be constant throughout the experiments. The mobility of all UEs was generated using the Simulation of Urban Mobility tool (SUMO) [51]. The ToD-UEs are moving continuously with a variable speed in the range of $(0,14]$ m/s corresponding to $(0,50]$ km/h, while the NonToD-UEs follow randomly generated trajectories inside the simulated topology, as provided by SUMO. We consider 10 simulated experiments, differing in the total background traffic load (depending on the number of NonToD-UEs) and the mobility patterns of the UEs. Table II summarizes the NS-3 parameters used in the simulated scenario.
@@
Encoder Complexity (MFLOPs) 90.5 90.5
Total Complexity (MFLOPs) 906 816
@@
To showcase the distribution of lower complexity computations, Table IV also includes the complexity of the total DISTINQT architecture (Total Complexity) along with the complexity of each NN-Encoder (Encoder Complexity) in MFLOPs (Mega Floating Point Operations)[52]. As it can be deduced from the table, the complexity of each NN-Encoder is approximately 10 times lower than that of the entire architecture. The maximum number of epochs for the learning phase was set to 1000. A min-max scaling method was applied to rescale the initial raw input data in the range $[0,1]$. Additionally, in order to avoid overfitting and potential degradation in the performance of the model, an early stopping function is applied, which monitors the Mean Squared Error (MSE) of the validation set [53]. As a result, the learning process stops, when the MSE score does not improve or starts deteriorating.
@@ V-A DISTINQT vs Centralized
Although centralized learning solutions present practical limitations (Section I), they are considered as high performing architectures in a variety of problems and use cases. As such, we compare the performance of our DISTINQT framework with a centralized LSTM-based approach, introduced in our previous work [55]. Table V, presents the evaluation results as average values of the MAE score for both input feature configurations and for all evaluation scenarios (50 in total). Two types of evaluation metrics are considered: (a) The ”overall MAE” score, representing the average MAE score over all prediction horizon timesteps, and (b) the ”20th Second MAE” score, representing the average MAE score of the last prediction horizon timestep only. The results under both the centralized, [55], and our proposed DISTINQT approaches are statistically very similar, with the DISTINQT showcasing a marginally better performance of 7.75% on average for both configurations and types of scores considered. Also c2 improves upon c1 on average by 5.65% for both scores, highlighting the importance of the additional input features of the MEC Server NET. MAE scores capture the overall ability of the model to yield a relatively low prediction error under the DISTINQT and the centralized approaches, considering typical uplink throughput values of the ToD applications (10-50 Mbps), as indicated by the EU funded 5GCroCo project [17].
@@ V-B Baseline Comparisons
V-C Data Privacy Preservation
To evaluate the effectiveness of the proposed data privacy methods discussed in Section III-C, an optimization problem is examined with the objective of approximating the initial raw input data. We consider the simplest network environment, comprised of one ToD-UE and one BS worker ($e\in$ToD-UE, BS}, $k=1$) using the input features of c1. A privacy violator attempts to approximate the raw input data sequences $\textbf{x}_{t}^{e,1}$, $\forall e$, assuming: 1) access to both ToD-UE and BS NN-Encoders, 2) knowledge of the exchanged context vectors, and the shape of the initial raw input data sequences, and 3) knowledge of the data normalization scale.
The privacy violator, using a uniform distribution, initializes the estimations $\hat{\textbf{x}}_{t}^{e,1}$, $\forall e$, in the range $[0,1]$ and uses these estimations along with the respective NN-Encoders to produce the context vectors $\hat{\textbf{v}_{t}}^{(e,1)}$. The objective is to minimize the MSE $d_{t}^{e,1}$ between the $\hat{\textbf{v}_{t}}^{(e,1)}$ and the original ones ($\textbf{v}_{t}^{(e,1)}$), defined as:
$\min_{\hat{\textbf{x}}_{t}^{e,1}}d_{t}^{e,1}=||\textbf{v}_{t}^{(e,1)}-\hat{\textbf{v}_{t}}^{(e,1)}||^{2},\forall e.\vspace{-5pt}$ (2)
[…]
$S_{e}={\frac{1}{1+u_{e}}}\cdot 100\%,\forall e,\vspace{-5pt}$ (3)
where
$u_{e}=||\textbf{x}_{t}^{e,1}-\hat{\textbf{x}}_{t}^{e,1}||_{2}.$ (4)
[…]
Figure 10: Original vs Estimated Sensitive Input Sequences
The overall similarity percentages for all features and for both NETs are approximately equal to $S_{\text{ToD-UE}}=18.69\%$ and $S_{\text{BS}}=23.17\%$.
It is important to note that, based on the results presented above, the estimation of the normalized input sequence is a challenging task. The lack of knowledge regarding the normalization method employed and the value range of each feature in each NET makes it impossible to obtain an initial raw input sequence that can be considered accurate.
@@ VI Conclusions
This paper introduces and demonstrate the effectiveness of DISTINQT, a novel privacy-aware distributed learning framework for QoS prediction in the presence of heterogeneous data types and model architectures. DISTINQT is based on sequence-to-sequence autoencoders, distributing various NN-Encoders across the network and contributing to data privacy preservation. Our framework is evaluated in the Tele-Operated Driving use case, where evaluation results show that it can reach and even surpass the performance of its centralized counterpart, while proving the validity of the privacy preserving claims. This is especially important for such safety critical use cases that require immediate and accurate decision making without compromising any sensitive information. Finally, DISTINQT achieves an average decrease in prediction error up to 65%, compared to six state-of-the-art centralized baseline solutions. Our future work will focus on finding a balance between the communication overhead that is introduced in the learning process and the robustness and generalization capabilities of the model.